Solve for "x"
logbase2 4 + logbase2 27= logbase2 x

Found 2 solutions by Alan3354, nerdybill:
Answer by Alan3354(21583)

Solve for "x"
logbase2 4 + logbase2 27= logbase2 x
--------------------
The base is the same, so that's not an issue.
The sum of 2 logs is the log of the product, so
log(4) + log(27) = log(4*27)
log(108) = log(x)
x = 108

Answer by nerdybill(5404)

We can write "logbase2 4" as "log2(4)".
.
You will need to apply "log rules". You can review them at:
http://www.purplemath.com/modules/logrules.htm
.
log2(4) + log2(27)= log2(x)
log2(4*27)= log2(x)
log2(108)= log2(x)
108 = x